# Unified Physical-Digital Attack Detection Challenge

### v1

@@ 2.1 Face Anti-spoofing datasets for Challenges
Print-Attack [1] addresses the prevalent issue of bypassing 2D FR systems using spoofed photographs. It comprises 400 samples: 200 genuine accesses and 200 videos using printed photos across 50 identities. Replay-Attack [3] is a dedicated resource focusing on two-dimensional PAs, which includes 1,300 video clips featuring various attack types with 50 subjects, recorded under controlled and adverse lighting conditions.. OULU-NPU [2] is a public PAD database designed to test the generalization of FR systems across different environments like lighting and background, various smartphones, and presentation attack instruments. It includes 5,940 high-resolution videos of 55 subjects, recorded in three environments with six smartphone models.
@@ 3.2 Challenge Protocol and Data Statistics
To thoroughly evaluate the performance of UAD frameworks, we established two distinct protocols within the UniAttackData framework. According to Tab. 2, Protocol 1 is designed to scrutinize performance across unified attack tasks. Protocol 2, on the other hand, is tailored to assess algorithmic generalization across ”unseen” attack types. Employing a ”leave-one-type-out testing” strategy, we further divide Protocol 2 into two sub-protocols, which enables a thorough investigation into the generalization capacity of FAS mechanisms against a broad and evolving threat landscape.
@@ 4.1 MTFace
Figure 1: The overall pipeline of MTFace’s method. The moire augmentation improves the model’s PAD ability, while self-blending augmentation improves the model’s DAD ability.
During pre-processing, all uncropped images are face-detected and cropped 20 pixels outward by the bound box. Finally, all live data in the training set is parsed to obtain the face mask area for further use in subsequent self-blending methods. As shown in Fig. 1, a moire data augmentation and a self-blending augmentation are applied to the different subsets based on the diverse structure of protocols. MTFace analyzes that screen replay will cause moire patterns on the surface texture of the image. Thus, the moire data augmentation method is designed and applied to protocol p1 and p2.1 so that the characteristics of physical attacks are introduced, improving the model’s ability to detect physical attacks across domains. Second, inspired by the work [43], a self-blending augmentation method is defined. For protocols p1 and p2.2, extra live-based data has been added with digital attack features by color, spatial transformation, and mask deformation. To achieve a more balanced training approach in response to the varying ratios of bona and spoof samples across three protocols, MTFace adjusted the cross-entropy loss weights accordingly. For protocol p1, they maintained an equal loss weight ratio of live to fake samples at 1:1, aiming for an even-handed learning process. In protocol p2.1, they adopted a 5:1 loss weight ratio, favoring live data to encourage the model to learn live features more effectively. For protocol p2.2, they applied a 2:1 loss weight ratio to better equilibrate the impact of different sample types on the model’s learning. This led to a marked enhancement in their experimental outcomes. Finally, ResNet-50 is set as the backbone, and the ImageNet pre-train weights are loaded. Their promising experiment results have beaten all other competitors in this challenge.
@@ 4.2 SeaRecluse
The team SeaRecluse presented a solution titled ”Cross-domain Face Anti-spoofing in Unified Physical-Digital Attack Dataset.” The team’s approach included several key steps and techniques: SCRFD is first applied to uncropped images in the training set for face cropping. For three protocols, the dataset partitioning methods differ: $80\%$ of the p1, $60\%$ of p2.1, and $40\%$ of p2.2’s training data are used along with the validation data for training, while the remaining data serves as the validation set. Regarding data augmentation and preprocessing, cropped images undergo no additional processing, whereas uncropped images are supplemented with data through loose and tight facial cropping. For each of the three tasks, distinct data augmentation measures are taken based on their respective focuses: Task P1 does not perform any enhancement operations; for P2.1, to balance the ratio of real face samples, real face data is downsampled and edge blank pixels are filled to reach a fixed size, effectively tripling the amount of real face data; in contrast, for P2.2, similar augmentation strategies are adopted for fake face data, introducing 4x and 8x downsampling, thereby increasing the fake face data volume by sevenfold. During the data review stage, some images’ incorrect aspect ratios were rectified and restored to their original proportions. All tasks implement normal data augmentation operations, with Task P2.1 also incorporating Gaussian blur. SeaRecluse chose ConvNeXt V2 backbone considering the competition requirements and local training resources for training. During the training phase, Image CutMix and label smoothing [36] techniques were leveraged to enhance the model’s generalization capability.
@@
The backbone model, the Swin-Large model [35], extracts features resulting in 1536-dimensional vectors, which, coupled with the image’s dimensional attributes, aid in differentiating between live and spoof images. Both pre-training and fine-tuning stages follow identical configurations across all protocols. This self-contained strategy yields a competitive edge in cross-attack-type scenarios and simplifies the transplanting of the method across various baselines. The simMIM strategy [50] is utilized with unlabeled training and development sets at the pre-training stage, involving segmenting an image into non-overlapping patches, concealing portions and prompting the model to infer the complete image. This masked image modeling (MIM) methodology fosters the model’s proficiency in feature extraction from incomplete visual information, which is helpful for protocol 2 in the testing phase. During the fine-tuning stage, they introduce a tailored data augmentation sequence. The Gaussian Noise augmentation represents digital threats, while ColorJitter and screen display simulations, including moire pattern overlays and gamma correction, are used to mimic physical attack types. These sequential augmentations are applied exclusively to training samples with designated probabilities.
@@ 4.4 BSP-Idiap
The Dual Branch Pixel-wise Binary Supervision (DBPixBiS) algorithm, presented by the BSP-Idiap team, is a novel FAS solution that extends their previous work DeepPixBiS. As shown in Fig. 3, their DBPixBiS utilizes a dual-branch network architecture that takes Fourier transform representations as input to capture both spatial and frequency domain properties of images. In the RGB branch of the model, central difference convolutional blocks are employed to process the visual information differently from standard convolutional operations. A vital aspect of the method lies in its pixel-wise binary supervision, where each pixel in the feature maps is guided by an attentional angular margin loss function during the training phase. This supervision strategy regularizes the learning process, reducing the over-fitting problem and promoting better generalization across diverse spoofing scenarios. The architecture includes an additional branch tailored for Fourier-transformed inputs to detect spoofing artifacts in the frequency domain. During inference, the mean value of the resulting feature map is used as the final spoofing detection score.
@@ 4.9 Cloud Recesses
The Cloud Recesses team has advanced a technique titled ”Random Masking for Face Anti-Spoofing Detection.” They proposed an innovative scheme, depicted in Fig. 6. The central premise of their method pivots on data augmentation through the deliberate obscuration of critical facial features. Initially, face landmarks are detected to inform the random masking process, intentionally veiling key areas such as the eyes, nose, and mouth. Subsequently, these modified images are channeled into the EfficientNet architecture for model training, targeting the enhancement of the model’s predictive acumen on validation sets. RetinaFace is employed for face detection in data pre-processing, yielding a 256x256 cropped face image. After that, dlib is utilized to detect 68 facial landmarks, outlining essential face components. During training, random occlusion of 3-5 regions per image is executed, attenuating the impact of identifiable facial features on the training outcome.
@@ 4.12 Inria-CENATAV-Tec
The team Inria-CENATAV-Tec refined UAD with their contribution, ”MobileNetV3-spoof with hyperparameter tuning.” Their methodology is encapsulated in Fig. 9, which details a flowchart of their pre-processing and UAD approach. In the pre-processing phase, the team employs ResNet-50 for landmark detection to ascertain face positioning. If the face is detected, it is aligned using the insightface template; otherwise, the image is resized and saved in its original state. The team utilizes the MobileNetV3-large-1.25 backbone for feature extraction and spoof detection, aiming to balance model complexity with performance. The model training leverages a Stochastic Gradient Descent optimizer with a multi-step learning rate strategy. Additionally, various augmentation techniques are applied to the data before normalization, aligning with each protocol’s mean and standard deviation. Their work aligns with the contemporary need for lightweight models that do not compromise the precision of FAS measures.
@@ 4.13 Vicognit
The team Vicognit presents ”FASTormer: Leveraging Vision Transformers for Face Anti-Spoofing,” a novel approach utilizing the Vision Transformer architecture for FAS. Their strategy is visualized in Fig. 10, which outlines the architecture of the proposed method. During the training phase, the team meticulously tunes the hyperparameters, such as learning rate and weight decay, to ensure that the model’s parameters are optimized effectively, leading to robust convergence and generalization. The Vicognit team’s solution hinges on a transformer-based model, indicating their focus on harnessing transformers’ capabilities to capture the intricate patterns necessary for distinguishing between bona and spoof face representations without reducing the dimensionality of the input data.

### v2

@@ 2.1 Face Anti-spoofing datasets for Challenges
Print-Attack [1] addresses the prevalent issue of bypassing 2D FR systems using spoofed photographs. It comprises 400 samples: 200 genuine accesses and 200 videos using printed photos across 50 identities. Replay-Attack [3] is a dedicated resource focusing on two-dimensional PAs, which includes 1,300 video clips featuring various attack types with 50 subjects, recorded under controlled and adverse lighting conditions. OULU-NPU [2] is a public PAD database designed to test the generalization of FR systems across different environments like lighting and background, various smartphones, and presentation attack instruments. It includes 5,940 high-resolution videos of 55 subjects, recorded in three environments with six smartphone models.
@@ 3.2 Challenge Protocol and Data Statistics
To thoroughly evaluate the performance of UAD frameworks, we established two distinct protocols within the UniAttackData framework. According to Tab. 2, Protocol 1 is designed to scrutinize performance across unified attack tasks. Protocol 2, on the other hand, is tailored to assess algorithmic generalization across "unseen" attack types. Employing a "leave-one-type-out testing" strategy, we further divide Protocol 2 into two sub-protocols, which enables a thorough investigation into the generalization capacity of FAS mechanisms against a broad and evolving threat landscape.
@@ 4.1 MTFace
Figure 1: The overall pipeline of MTFace’s method. The moiré augmentation improves the model’s PAD ability, while self-blending augmentation improves the model’s DAD ability.
During pre-processing, all uncropped images are face-detected and cropped 20 pixels outward by the bound box. Finally, all live data in the training set is parsed to obtain the face mask area for further use in subsequent self-blending methods. As shown in Fig. 1, a moiré data augmentation and a self-blending augmentation are applied to the different subsets based on the diverse structure of protocols. MTFace analyzes that screen replay will cause moiré patterns on the surface texture of the image. Thus, the moiré data augmentation method is designed and applied to protocol p1 and p2.1 so that the characteristics of physical attacks are introduced, improving the model’s ability to detect physical attacks across domains. Second, inspired by the work [45], a self-blending augmentation method is defined. For protocols p1 and p2.2, extra live-based data has been added with digital attack features by color, spatial transformation, and mask deformation. To achieve a more balanced training approach in response to the varying ratios of bona and spoof samples across three protocols, MTFace adjusted the cross-entropy loss weights accordingly. For protocol p1, they maintained an equal loss weight ratio of live to fake samples at 1:1, aiming for an even-handed learning process. In protocol p2.1, they adopted a 5:1 loss weight ratio, favoring live data to encourage the model to learn live features more effectively. For protocol p2.2, they applied a 2:1 loss weight ratio to better equilibrate the impact of different sample types on the model’s learning. This led to a marked enhancement in their experimental outcomes. Finally, ResNet-50 is set as the backbone, and the ImageNet pre-train weights are loaded. Their promising experiment results have beaten all other competitors in this challenge.
@@ 4.2 SeaRecluse
The team SeaRecluse presented a solution titled "Cross-domain Face Anti-spoofing in Unified Physical-Digital Attack Dataset." The team’s approach included several key steps and techniques: SCRFD is first applied to uncropped images in the training set for face cropping. For three protocols, the dataset partitioning methods differ: $80\%$ of the p1, $60\%$ of p2.1, and $40\%$ of p2.2’s training data are used along with the validation data for training, while the remaining data serves as the validation set. Regarding data augmentation and preprocessing, cropped images undergo no additional processing, whereas uncropped images are supplemented with data through loose and tight facial cropping. For each of the three tasks, distinct data augmentation measures are taken based on their respective focuses: Task P1 does not perform any enhancement operations; for P2.1, to balance the ratio of real face samples, real face data is downsampled and edge blank pixels are filled to reach a fixed size, effectively tripling the amount of real face data; in contrast, for P2.2, similar augmentation strategies are adopted for fake face data, introducing 4x and 8x downsampling, thereby increasing the fake face data volume by sevenfold. During the data review stage, some images’ incorrect aspect ratios were rectified and restored to their original proportions. All tasks implement normal data augmentation operations, with Task P2.1 also incorporating Gaussian blur. SeaRecluse chose ConvNeXt V2 backbone considering the competition requirements and local training resources for training. During the training phase, Image CutMix and label smoothing [38] techniques were leveraged to enhance the model’s generalization capability.
@@
The backbone model, the Swin-Large model [37], extracts features resulting in 1536-dimensional vectors, which, coupled with the image’s dimensional attributes, aid in differentiating between live and spoof images. Both pre-training and fine-tuning stages follow identical configurations across all protocols. This self-contained strategy yields a competitive edge in cross-attack-type scenarios and simplifies the transplanting of the method across various baselines. The simMIM strategy [52] is utilized with unlabeled training and development sets at the pre-training stage, involving segmenting an image into non-overlapping patches, concealing portions and prompting the model to infer the complete image. This masked image modeling (MIM) methodology fosters the model’s proficiency in feature extraction from incomplete visual information, which is helpful for protocol 2 in the testing phase. During the fine-tuning stage, they introduce a tailored data augmentation sequence. The Gaussian Noise augmentation represents digital threats, while ColorJitter and screen display simulations, including moiré pattern overlays and gamma correction, are used to mimic physical attack types. These sequential augmentations are applied exclusively to training samples with designated probabilities.
@@ 4.4 BSP-Idiap
The Dual Branch Pixel-wise Binary Supervision (DBPixBiS) algorithm, presented by the BSP-Idiap team, is a novel FAS solution that extends their previous work DeepPixBiS [10]. As shown in Fig. 3, their DBPixBiS utilizes a dual-branch network architecture that takes Fourier transform representations as input to capture both spatial and frequency domain properties of images. In the RGB branch of the model, central difference convolutional blocks [54] are employed to process the visual information differently from standard convolutional operations. A vital aspect of the method lies in its pixel-wise binary supervision, where each pixel in the feature maps is guided by an attentional angular margin loss [14] function during the training phase. This supervision strategy regularizes the learning process, reducing the over-fitting problem and promoting better generalization across diverse spoofing scenarios. The architecture includes an additional branch tailored for Fourier-transformed inputs to detect spoofing artifacts in the frequency domain. During inference, the mean value of the resulting feature map is used as the final spoofing detection score.
@@ 4.9 Cloud Recesses
The Cloud Recesses team has advanced a technique titled "Random Masking for Face Anti-Spoofing Detection." They proposed an innovative scheme, depicted in Fig. 6. The central premise of their method pivots on data augmentation through the deliberate obscuration of critical facial features. Initially, face landmarks are detected to inform the random masking process, intentionally veiling key areas such as the eyes, nose, and mouth. Subsequently, these modified images are channeled into the EfficientNet architecture for model training, targeting the enhancement of the model’s predictive acumen on validation sets. RetinaFace is employed for face detection in data pre-processing, yielding a 256x256 cropped face image. After that, dlib is utilized to detect 68 facial landmarks, outlining essential face components. During training, random occlusion of 3-5 regions per image is executed, attenuating the impact of identifiable facial features on the training outcome.
@@ 4.12 Inria-CENATAV-Tec
The team Inria-CENATAV-Tec refined UAD with their contribution, "MobileNetV3-spoof with hyperparameter tuning." Their methodology is encapsulated in Fig. 9, which details a flowchart of their pre-processing and UAD approach. In the pre-processing phase, the team employs ResNet-50 for landmark detection to ascertain face positioning. If the face is detected, it is aligned using the insightface template; otherwise, the image is resized and saved in its original state. The team utilizes the MobileNetV3-large-1.25 backbone for feature extraction and spoof detection, aiming to balance model complexity with performance. The model training leverages a Stochastic Gradient Descent optimizer with a multi-step learning rate strategy. Additionally, various augmentation techniques are applied to the data before normalization, aligning with each protocol’s mean and standard deviation. Their work aligns with the contemporary need for lightweight models that do not compromise the precision of FAS measures.
@@ 4.13 Vicognit
The team Vicognit presents "FASTormer: Leveraging Vision Transformers for Face Anti-Spoofing," a novel approach utilizing the Vision Transformer architecture for FAS. Their strategy is visualized in Fig. 10, which outlines the architecture of the proposed method. During the training phase, the team meticulously tunes the hyperparameters, such as learning rate and weight decay, to ensure that the model’s parameters are optimized effectively, leading to robust convergence and generalization. The Vicognit team’s solution hinges on a transformer-based model, indicating their focus on harnessing transformers’ capabilities to capture the intricate patterns necessary for distinguishing between bona and spoof face representations without reducing the dimensionality of the input data.